3x+2y=11
7x-y=3
Need ti know how to solve this problem

Answer by Maths68(1474)

3x+2y=11...........(1)
7x-y=3.............(2)
Multiply (2) by 2 and add to (1)
3x+2y=11
14x-2y=6
-----------------
17x=17
Divide by 17 both sides
17x/17=17/17
x=1
Put the value of x in (1)
3x+2y=11...........(1)
3(1)+2y=11
3+2y=11
2y=11-3
2y=8
Divide by 2 both sides
2y/2=8/2
y=4
x,y=1,4


Check
=======
Put the values of x and y in (1)
3x+2y=11...........(1)
3(1)+2(4)=11
3+8=11
11=11
Put the values of x and y in (2)
7x-y=3.............(2)
7(1)-(4)=3
7-4=3
3=3